(g)+ F2 g) Estinate the carbon-fluorine bood energy given that the CC bond energy is 347 kJ/mol,the C-C bond energy is 614 kJmol, and the P-P bond energy is 154 kJ/mol. Cartoo Borine bodo kJ/mol 10 iten attempts remaining
Answer #1

Heat required to dissociate C2H4 and F 2 is

For breaking                           1 C=C 614 KJ / mol

For breaking 4 C----H                             2 x 413 KJ / mol

For breaking F-----F 154 KJ / mol

                              Total = 614 + ( 4 x 413 ) + 154 =2420 KJ / mol

If same atoms combine to give C2H2F2 energy is given out

For making                          1 C-------C                           347 KJ / mol

For making                          4 C-----H                             4 x 413 KJ / mol

For making                          2 C------F 2 x Z KJ / mol

Here Z is bond dissociation energy of C----Cl bond

                                 Total = 347 + (4 x 413) + (2 x Z) = - (1999+( 2 x Z ) ) KJ/ mol

\DeltaH reaction = - (1999 + (2 x Z)) KJ/ mol + 2420 KJ / mol = -549 KJ / mol

                                                        (1999 + (2 x Z)) KJ/ mol = -549 KJ / mol -   2420 KJ / mol

                                                        -1999- (2 x Z)) KJ/ mol = -2969 KJ / mol

                                                               - (2 x Z)) KJ/ mol = -2677 KJ / mol +1999

                                                                                            = -970 KJ/ mol

                                                                                         Z = (970 KJ/ mol)/2

                                                                                         Z = 485 KJ/mol

Carbon fluorine bond energy = 485 KJ / mol